OpenStudy (anonymous):

Which answer best describes the solution to the equation 5z + 4 = 4 + 5z? A. one solution B. no solution C. an infinite number of solutions D. z = 20

OpenStudy (anonymous):

@abb0t

OpenStudy (nincompoop):

z=z

OpenStudy (nincompoop):

identity property

Nnesha (nnesha):

subtract 4 both side and 5z

Nnesha (nnesha):

4 and 5z should cancel out then you have left 0 =0 which mean ??? not one solution

OpenStudy (nincompoop):

there is an infinite number of solution that is any value can be Z

OpenStudy (anonymous):

so it would be B

OpenStudy (nincompoop):

uhhh don't listen to her

Nnesha (nnesha):

nope when you get both side equal 0=0 or -6=-6 its mean an infinite number of solutions

OpenStudy (nincompoop):

no solution would mean you didn't reach an equality

Nnesha (nnesha):

no solution when 4 = 8 when both side are not equal

Nnesha (nnesha):

hope thats help :)

OpenStudy (anonymous):

oh so it would be C then yes @nincompoop @Nnesha

Nnesha (nnesha):

yesh

OpenStudy (nincompoop):

whenever you see an identity z = z z+1 = z+1 50z + 10 = 10 + 50z they're all offer you an infinite set of solution meaning that any value can replace z and you would still reach an equality

OpenStudy (anonymous):

thanks guys i got any 80% :D